# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

### v2

@@ 3.2.4 Training and Inference
During the training process, our model extracts single event at a time. Our training loss consists of two parts: classification loss $\mathcal{L}_{c}$ and boundary loss $\mathcal{L}_{b}$, and $\lambda$ is hyperparameter.
$\mathcal{L}=\mathcal{L}_{c}+\lambda\mathcal{L}_{b}.$ (10)
@@
Figure 4: The main structure of CARLG involves an input sequence, encompassing roles and event types, passing through a transformer-based encoder. This encoder generates output in the form of context representations, role representations, and attention heads. The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations. Subsequently, the context vector is combined with the latent role vector to create the ultimate span representation. Finally, the classification module leverages this representation to forecast argument roles for all potential spans.
4 CARLG Model
In this section, we introduce Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles. These two tecniques are then well equipped on our proposed baseline and our CARLG model are shown in Figure 4.
4.1 Contextual Clues Aggregation
The average pooling representation that previous span-based methods xu2022two ; ebner2020multi ; zhang2020two disregards important contextual information from non-argument words. Although the self-attention mechanism of the pre-trained encoder can capture token-level interactions, it lacks specific global interaction tailored to the event and candidate arguments. To address this limitation, we propose a Contextual Clues Aggregation (CCA) module that selects and combines relevant contextual information for each tuple comprising a candidate span and the event trigger word, denoted as $(s_{i:j},t)$. In this module, as shown in Figure 5, we leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers.
We employ the token-level attention heads $\textbf{A}^{w}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model. Here, $H$ represents the number of attention heads, and $l_{w}$ signifies the length of word pieces. Utilizing these attention heads, we compute the context attention $\textbf{A}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ for each candidate span spanning from $w_{i}$ to $w_{j}$, by applying average pooling:
$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{w}_{h,m}.$ (11)
Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}^{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
Figure 5: Contextual Clues Aggregation Module.
$\begin{array}[]{c}\textbf{p}^{c}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}^{s_{i:j}}={\mathbf{H}^{w}}^{T}\ \textbf{p}^{c}_{i:j},\end{array}$ (12)
We can apply our CCA module to other span-based methods by modifying Eq. 2 and fusing the contextual clue information $\textbf{c}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{c}_{i:j}=tanh(\mathbf{W}^{CCA}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}^{s_{i:j}}]),$ (13)
where $\mathbf{W}^{CCA}\in\mathbb{R}^{2d\times d}$ is learnable parameter. $\textbf{s}^{c}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
Figure 6: Role-based Latent Information Guidance Module. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
4.2 Role-based Latent Information Guidance
In this section, we propose Role-based Latent Information Guidance (RLIG) module, which is shown in Figure 6.
4.2.1 Role-interactive Encoding
To capture the semantic correlations between roles effectively, we integrate the information of role types into the input and utilize multi-head attention to facilitate interaction between the context and roles. This enables us to obtain the context and role representation in a shared knowledge space. We generate latent role embeddings by assigning unique special tokens   [^2] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 to embody different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation. Considering that role names carry valuable semantic information wangetal2022query , we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation. This approach ensures the capture of unique attributes and subtle semantic distinctions of each role. For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
Given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input sequence described in Sec. 3.2.1 by concatenating these elements in the following manner:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]}$
$[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$,
where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the ”none” category. By employing role-interactive encoding, we are able to get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start event $[E_{e}]$, the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{r}\in\mathbb{R}^{l_{r}\times d}$, all from $\mathbf{H}^{s}$. Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
4.2.2 Role Information Fusion
Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context. To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information. The aggregated role information $\textbf{r}^{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Equations 11 and 12:
$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{r}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{r}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}^{s_{i:j}}={\mathbf{H}^{r}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (14)
where $\textbf{A}^{r}\in\mathbb{R}^{H\times l_{w}\times l_{r}}$ represents the role attention heads obtained from the final transformer layer of the pre-trained model. $\textbf{A}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ represents the role attention for each candidate span, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$. Additionally, the vector $\textbf{p}^{r}_{i:j}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles. We can modify Eq. 13 and integrate the latent role information $\textbf{r}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{r}_{i:j}=tanh(\mathbf{W}^{RLIG}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{r}^{s_{i:j}}]),$ (15)
where $\mathbf{W}^{RLIG}\in\mathbb{R}^{2d\times d}$ is learnable parameter.
4.2.3 Latent Role Information Guidance
We first map the original $d$-dimensional role features into a reduced dimension $d^{\prime}$ via a fully connected network to obtain the latent role information $\mathbf{H}^{r^{\prime}}\in\mathbb{R}^{l_{r}\times d^{{}^{\prime}}}$ as follows:
$\mathbf{H}^{r^{\prime}}=\mathbf{H}^{r}\mathbf{W}_{3}+\mathbf{b}_{w},$ (16)
where $\mathbf{W}_{3}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$ and $\mathbf{b}_{w}\in\mathbb{R}^{d^{{}^{\prime}}}$ are learnable parameters. To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  tucker1964extension . We define candidate span-trigger-role triples set $\mathcal{T}=\{<s_{i:j},r_{k},t_{e}>|s_{i:j}\in\mathcal{S},r_{k}\in\mathcal{R}_{e}\}$. Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$ calculated in Eq. 7, and for simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
where ${\widehat{\mathbf{r}}}_{\mathrm{k}}=\mathbf{H}^{r^{\prime}}_{k,:}$ represents the transformed role features and ×$\times_{n}$ denotes the tensor product operation along the $n^{th}$ dimension. The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference 8237586 .
@@ 5.1.1 Datasets
We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS ebner2020multi , WikiEvents lietal2021document and MLEE 10.1093/bioinformatics/bts407 , and detailed data statistics are listed in Table 1.
@@
  • 1.
    RAMS  The dataset includes 9,124 annotated events sourced from English online news articles. In this dataset, each event forms a single instance, and events that occur within the same context are grouped together into a single instance containing multiple events. We adopt the original train/dev/test split following ebner2020multi .
  • 2.
@@
  • 3.
    MLEE  This dataset is a document-level event extraction dataset that focuses on bio-medical publications, consisting of manually annotated abstracts written in English. We follow the procedure outlined in 10.1093/bioinformatics/btaa540 and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set. In this case, we use the training set as the development set for our experiments following  10.1093/bioinformatics/btaa540 .
@@ 5.1.2 Evaluation Metrics
For the RAMS dataset, we adopt the method detailed in xu2022two , focusing on both Span F1 and Head F1 scores across the development and test sets. Span F1 evaluates the precision of matching predicted argument spans against the gold standard, while Head F1 quantifies the accuracy of identifying the central word in each argument span. For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by ma-etal-2022-prompt . Regarding the the WikiEvents dataset, we present both Head F1 and Coref F1 metrics on the test set for Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, consistent with the methodology described in lietal2021document . The Coref F1 score, inspired by the evaluation technique from jigrishman2008refining , gauges the coreference accuracy between extracted arguments and the established gold standard. For the MLEE dataset, we report Span F1 scores on the test set for both Arg IF and Arg CF tasks, following the methodologies in lietal2021document ; ma-etal-2022-prompt .
@@ 5.1.3 Baselines
Table 2: Detailed hyperparameter settings.
@@
Alpha 10 10 10
Gamma 2 2 2
@@ 5.1.4 Hyperparameter Setting
Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of TSAR xu2022two . More detailed hyperparameters settings are listed in Table 2.
Table 3: The main results for the RAMS dataset. Arg IF denotes the F1 score for the Argument Identification subtask, and other metrics are for the Argument Classification subtask. * indicates results reported by  ma-etal-2022-prompt . The highest scores are highlighted in bold, while the second-highest scores are underlined.
@@
Tag-based
BERT-CRF ebner2020multi 38.1 45.7 39.3 47.1 -
@@
Generation-based
FEAE weietal2021trigger - - 47.40 - -

### v4

@@ 5.1.4 Training and Inference
Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss (Lin et al., 2017) during the training process. Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(\varphi(s_{i:j},r_{k},t_{e})\;=y_{s_{i:j}}}))]}^{\gamma}\vspace{1.3ex}\\ \cdot\log P(\varphi(s_{i:j},r_{k},t_{e})\;=y_{s_{i:j}}).\end{array}$ (10)
@@
5.2 $\text{CARLG}_{\text{prompt}}$
In this section, we apply our CARLG framework to prompt-based EAE approaches, where the architecture is shown in Figure 7. We choose the architecture of PAIE (Ma et al., 2022b) as the foundation to illustrate the prompt-based baseline. Here, we only provide a detailed introduction to the parts related to our CCA and RLIG and the other parts are omitted.
@@ 5.2 $\text{CARLG}_{\text{prompt}}$
Figure 6: The main structure of $\text{CARLG}_{\text{prompt}}$ involves an input sequence, encompassing roles and prompts, passing through an encoder-decoder architecture. This encoder generates output in the form of context representations, role representations, and attention heads. The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations. Subsequently, the context vector and the latent role vector are sent to the span selector to predict corresponding spans of each slot. Ultimately, based on the predicted spans, we extract the corresponding argument for each role from the text.
5.2.1 Input Preprocess
We utilize manual templates to generate the prompt $P$ for each instance [^3] Inspired by this work (Ma et al., 2022b), manual template has the most stable performance and usually the better result than other prompt variants.. The template for each event type is pre-defined as a sentence with slots $role\right>$. The generated prompt for each specific event type is a filled template where slots are substituted with actual role names. We use $s_{k}$ to denote the $k$-th slot in the prompt $P$.
[…]
The above case shows a concrete example for the prompt creation, and $s_{0}$ = $victim\right>$, $s_{1}$ = $place\right>$, $s_{2}$ = $killer\right>$.
Next, we concatenate the input text with the prompt. Following Algorithm 2, we prepend the role type to the sequence to obtain the final sequence $S_{\text{prompt}}$.
5.2.2 Prompt-based CCA and RLIG
We adopt an encoder-decoder architecture. The encoder is employed to encode the input text, while the decoder is tasked with deriving the event-oriented context and context-oriented prompt representation $\mathbf{H}_{\mathrm{de}}$:
$\begin{array}[]{c}[\mathbf{A};\mathbf{H}_{\mathrm{en}}]=\mathrm{Encoder}(S_{\text{prompt}}),\vspace{1.0ex}\end{array}$ (11)
where the $\mathrm{Encoder}$ is a pre-trained transformer-based encoder. $\mathbf{A}\in\mathbb{R}^{H\times l_{s}\times l_{s}}$ is the multi-head attention matrix and $\mathbf{H}_{\mathrm{en}}$. $H$ is the attention head numbers and $l_{s}$ is the length of input sequence $S_{\text{prompt}}$.
For prompt-based EAE approaches, the representation of event arguments is indicated through slots in the prompt. Therefore, we replace the $\textbf{A}^{C}_{a_{i}}$ in Algorithm 1 and $\textbf{A}^{R}_{a_{i}}$ in Algorithm 2 with the corresponding attention heads of the slots in the input sequence, respectively. The remaining operations are consistent with the steps in Algorithm 1 and Algorithm 2. Ultimately, we can obtain the enhanced context vector $\textbf{c}_{s_{i}}$ and role vector $\textbf{r}_{s_{i}}$ for the the slot of argument $a_{i}$.
Then we utilize a decoder to derive the event-oriented context and context-oriented prompt representation $\mathbf{H}_{\mathrm{de}}$:
$\begin{array}[]{c}\mathbf{H}_{\mathrm{de}}=\mathrm{Decoder}(\mathbf{H}_{\mathrm{en}}),\vspace{1.0ex}\end{array}$ (12)
where the $\mathrm{Decoder}$ is a transformer-based decoder and $\mathbf{H}_{\mathrm{de}}\in\mathbb{R}^{l_{s}\times d}$. Next, we integrate $\textbf{c}_{s_{i}}$ and $\textbf{r}_{s_{i}}$ into the decoder output $\mathbf{h}_{s{i}}\in\mathbb{R}^{d}$ for slot $s_{i}$, resulting in an updated representation $\tilde{\mathbf{h}}{s{i}}\in\mathbb{R}^{d}$::
$\tilde{\mathbf{h}}_{s_{i}}=\mathrm{tanh}(\mathbf{W}^{Prompt}[\mathbf{h}_{s_{i}};\textbf{c}_{s_{i}};\textbf{r}_{s_{i}}]),$ (13)
where $\mathbf{W}^{Prompt}\in\mathbb{R}^{3d\times d}$ is learnable parameter.
5.2.3 Span Selection
Upon acquiring the final representation $\tilde{\mathbf{h}}_{s{i}}$ for every slot across each event, we employ a methodology inspired by (Ma et al., 2022a) to convert these into a duo of span selectors $\Phi_{s_{i}}^{\mathrm{start}},\Phi_{s_{i}}^{\mathrm{end}}$:
$\begin{array}[]{c}\Phi_{s_{i}}^{\mathrm{start}}=\tilde{\mathbf{h}}_{s{i}}\circ\mathbf{w}_{\mathrm{start}},\vspace{1.0ex}\ \Phi_{s_{i}}^{\mathrm{end}}=\tilde{\mathbf{h}}_{s{i}}\circ\mathbf{w}_{\mathrm{end}},\end{array}$ (14)
where $\circ$ denotes element-wise multiplication and $\mathbf{w}_{\mathrm{start}},\mathbf{w}_{\mathrm{end}}\in\mathbb{R}^{d}$ are tunable parameters. These selectors, $\Phi_{s_{i}}^{\mathrm{start}}$ and $\Phi_{s_{i}}^{\mathrm{end}}$, are pivotal in identifying the start and end points of the slot $s_{i}$ within the original text. Following the approach in (Ma et al., 2022a), the Bipartite Matching Loss is utilized (Carion et al., 2020) to refine the alignment of predicted and actual argument spans during training ( For detailed content on this part, please refer to the work PAIE (Ma et al., 2022b), and we follow their approach here.)
@@ 6.1.1 Datasets
We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS (Ebner et al., 2020), WikiEvents (Li et al., 2021) and MLEE (Pyysalo et al., 2012), and detailed data statistics are listed in Table 1.
@@
  • RAMS  The dataset includes 9,124 annotated events sourced from English online news articles. In this dataset, each event forms a single instance, and events that occur within the same context are grouped together into a single instance containing multiple events. We adopt the original train/dev/test split following (Ebner et al., 2020).
@@
  • MLEE  This dataset is a document-level event extraction dataset that focuses on bio-medical publications, consisting of manually annotated abstracts written in English. We follow the procedure outlined in (Trieu et al., 2020) and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set. In this case, we use the training set as the development set for our experiments following  (Trieu et al., 2020).
@@ 6.1.2 Evaluation Metrics
For the RAMS dataset, we adopt the method detailed in (Xu et al., 2022b), focusing on both Span F1 and Head F1 scores across the development and test sets. Span F1 evaluates the precision of matching predicted argument spans against the gold standard, while Head F1 quantifies the accuracy of identifying the central word in each argument span. For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by (Ma et al., 2022b). Regarding the Wikievents and MLEE datasets, we report Span F1 scores on the test set for both Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, following the methodologies in (Li et al., 2021; Ma et al., 2022b).
@@ 6.1.3 Baselines
Table 2: Hyperparameter settings for $\text{CARLG}_{\text{span}}$. * means that we tuned the hyperparameters in our experiments. The rest of hyperparameters are set the same as TSAR (Xu et al., 2022b).
@@
Alpha* 10 10 10
Gamma* 2 2 2
Hyperparameters RAMS WikiEvents MLEE
Training Steps 10000 10000 10000
Warmup Ratio* 0.1 0.1 0.2
Learning Rate* 2e-5 3e-5 3e-5
Max Gradient Norm 5 5 5
Batch Size* 4 4 4
Context Window Size 250 250 250
Max Span Length 10 10 10
Max Encoder Seq Length 500 500 500
Max Prompt Length* 210 360 360
Encoder Layers* 17 17 17
Decoder Layers* 7 7 7
Table 3: Hyperparameter settings for $\text{CARLG}_{\text{prompt}}$. * means that we tuned the hyperparameters in our experiments. The rest of hyperparameters are set the same as PAIE (Ma et al., 2022a).
@@ 6.1.4 Hyperparameter Setting
Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of baselines. Here, we provide the hyperparameter settings for the span-based method TSAR in Table 2 and the prompt-based method PAIE in Table 3
Table 4: The main results for the RAMS dataset. Arg IF denotes the F1 score for the Argument Identification subtask, and other metrics are for the Argument Classification subtask. Bold indicates the best experimental results. The values in parentheses represent the performance compared to the original methods after incorporating our approach.
@@
Generation-based Methods
EEQA (Du & Cardie, 2020b) - - 46.7 - 48.7
FEAE (Wei et al., 2021) - - 47.4 - -
BART-Gen (Li et al., 2021) - - 48.6 57.3 51.2
${\text{EA}^{2}\text{E}}$ (Zeng et al., 2022) 46.1 53.9 47.8 56.4 53.2
EACE (Zhou et al., 2024) - - 50.1 56.3 -
Span-based Methods
Two-Step (Zhang et al., 2020d) 38.9 46.4 40.1 47.7 46.5
@@
TSAR (Xu et al., 2022b) 49.2 56.8 51.2 58.5 56.1
SCPRG (Liu et al., 2023b) 50.2 57.2 51.9 59.1 -
Prompt-based Methods
PAIE (Ma et al., 2022b) 50.9 56.5 52.2 58.8 56.8
RKDE (Hu et al., 2023) - - 50.3 55.1 -
HRA (Ren et al., 2023) - - 48.4 - 54.6
SPEAE (Nguyen, 2023) - - 52.5 - 57.0
TabEAE (He et al., 2023) 51.9 57.8 51.8 59.0 56.7
$\text{CARLG}_{\text{span}}$ (Ours)
Two-Step+$\text{CARLG}_{\text{span}}$ 43.4 ($\uparrow$4.5) 50.6 ($\uparrow$4.2) 45.3 ($\uparrow$5.2) 52.8 ($\uparrow$5.1) 50.3 ($\uparrow$3.8)
${\text{TSAR}}_{{\text{large}}}$+$\text{CARLG}_{\text{span}}$ 50.7 ($\uparrow$1.5) 58.1 ($\uparrow$1.3) 52.6 ($\uparrow$1.4) 59.9 ($\uparrow$1.4) 57.1 ($\uparrow$1.0)
$\text{CARLG}_{\text{prompt}}$ (Ours)
PAIE+$\text{CARLG}_{\text{prompt}}$ 52.8 ($\uparrow$1.9) 58.7 ($\uparrow$2.2) 54.4 ($\uparrow$2.2) 60.7 ($\uparrow$1.9) 59.1 ($\uparrow$2.3)
TabEAE+$\text{CARLG}_{\text{prompt}}$ 53.1 ($\uparrow$1.2) 59.2 ($\uparrow$1.4) 54.3 ($\uparrow$2.5) 60.3 ($\uparrow$1.3) 58.6 ($\uparrow$1.9)
6.2 Main Results
In this section, we compare the effectiveness of our CARLG framework on three datasets. We validate the CARLG framework both on span-based and prompt-based methods and then conduct further analyses.
6.2.1 Performance of Span-based Methods equipped with CARLG
As shown in Table 4 and Table 5, both Two-Step and TSAR exhibit improved performance when equipped with $\text{CARLG}_{\text{span}}$. Two-Step yields an improvement of +5.2/+5.1 Span F1/Head F1 for Arg CF and +3.8 Span F1 for Arg IF on RAMS, +6.9/+6.1 Span F1 for Arg IF/Arg CF on WikiEvents, and +2.2/+1.9 Span F1 for Arg IF/Arg CF on MLEE. TSAR achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.4/+1.4), Span F1 for Arg IF/Arg CF on WikiEvents (+1.5/+1.5), and Span F1 for Arg IF/Arg CF on MLEE (+1.3/+1.2). It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest. This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration. These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
6.2.2 Performance of Prompt-based Methods equipped with CARLG
When equipped with $\text{CARLG}_{\text{prompt}}$, as illustrated in Table 4 and Table 5, the prompt-based methods PAIE and TabEAE show significant enhancements across all three datasets, surpassing previous generation-based and span-based models. Specifically, PAIE shows a boost of +1.1/+1.4 in Span F1/Head F1 for Arg CF and +2.3 in Span F1 for Arg IF on RAMS, +2.1/+1.7 in Span F1 for Arg IF/Arg CF on WikiEvents, and +2.0/+1.3 in Span F1 for Arg IF/Arg CF on MLEE. Meanwhile, TabEAE achieves remarkable improvement in Span F1/Head F1 for Arg CF on RAMS (+2.5/+1.3), in Span F1 for Arg IF/Arg CF on WikiEvents (+0.7/+1.1), and in Span F1 for Arg IF/Arg CF on MLEE (+1.1/+0.9). These results validate the effective adaptability of our CARLG framework to prompt-based EAE methods, showcasing enhanced extraction capabilities. This is attributed to its adept utilization of contextual clues and the integration of latent role representations enriched with semantic connections.
Table 5: Main results of WikiEvents and MLEE datasets. Bold indicates the best experimental results. The values in parentheses represent the performance compared to the original methods after incorporating our approach.
WikiEvents MLEE
Method Arg IF Arg CF Arg IF Arg CF
Span F1 Span F1 Span F1 Span F1
Generation-based Methods
EEQA (Du & Cardie, 2020b) 56.9 54.5 70.3 68.7
${\text{EA}^{2}\text{E}}$ (Zeng et al., 2022) 68.8 63.7 70.6 69.1
EACE (Zhou et al., 2024) 71.1 66.2 - -
Span-based Methods
Two-Step (Zhang et al., 2020d) 52.9 51.1 66.7 61.9
[…]
TSAR (Xu et al., 2022b) 70.8 65.5 72.3 71.3
SCPRG (Liu et al., 2023b) 71.3 66.4 - -
Prompt-based Methods
PAIE (Ma et al., 2022b) 70.5 65.3 72.1 70.8
RKDE (Hu et al., 2023) 69.1 63.8 - -
HRA (Ren et al., 2023) 69.6 63.4 - -
SPEAE (Nguyen, 2023) 71.9 66.1 - -
TabEAE (He et al., 2023) 71.1 66.0 75.1 74.2
$\text{CARLG}_{\text{span}}$ (Ours)
Two-Step+$\text{CARLG}_{\text{span}}$ 59.8 ($\uparrow$6.9) 57.2 ($\uparrow$6.1) 68.9 ($\uparrow$2.2) 63.8 ($\uparrow$1.9)
${\text{TSAR}}_{{\text{large}}}$+$\text{CARLG}_{\text{span}}$ 72.1 ($\uparrow$1.5) 66.8 ($\uparrow$1.5) 73.4 ($\uparrow$1.3) 72.0 ($\uparrow$1.2)
$\text{CARLG}_{\text{prompt}}$ (Ours)
PAIE+$\text{CARLG}_{\text{prompt}}$ 72.6 ($\uparrow$2.1) 67.0 ($\uparrow$1.7) 74.1 ($\uparrow$2.0) 73.1 ($\uparrow$1.3)
TabEAE+$\text{CARLG}_{\text{prompt}}$ 72.1 ($\uparrow$1.0) 67.1 ($\uparrow$1.1) 76.2 ($\uparrow$1.1) 75.1 ($\uparrow$0.9)
6.2.3 Further analyses
Further analyzing the experimental results in Table 4 and Table 5, we observe that: (1) Our CARLG framework shows relatively larger improvement on simple models such as Two-step and PAIE. We hypothesize this may be because other components of complex models have synergistic or antagonistic effects on CCA and RLIG modules, constraining the enhancement potential of the CARLG modules and preventing them from fully unleashing their effectiveness. (2) The prompt-based methods yield the best results, with those equipped with the CARLG framework outperforming both generation-based and span-based approaches in terms of performance.
[…]
RAMS WikiEvents
Method Params Arg CF Arg IF Arg IF Arg CF
Span F1 Head F1 Span F1 Span F1 Span F1
$\text{PAIE}_{\text{C}}$ 408.22M 52.81±0.43 58.74±0.36 57.03±0.87 72.62±0.98 67.02±0.85
    -CCA 407.13M 51.28±0.60 56.67±0.51 55.43±0.48 70.16±0.87 65.87±1.06
    -RLIG 406.14M 51.77±0.36 56.96±0.44 55.91±0.48 70.83±0.79 66.49±0.68
    -both 405.12M 50.94±0.36 56.47±0.22 55.02±0.24 69.85±0.45 65.12±0.58
[…]
    -CCA 373.65M 40.24±0.84 47.66±0.67 46.98±0.26 55.48±1.87 53.65±1.02
    -RLIG 371.68M 41.67±1.12 48.26±0.95 47.24±0.60 56.43±1.46 54.87±1.19
    -both 370.63M 38.93±0.80 46.44±0.66 46.10±0.90 52.92±0.56 51.09±1.22
6.3 Ablation Study